OpenStudy (anonymous):

How do I prove that the function y=sqrt(2x+3) is only continuous past -1.5 with limits?

OpenStudy (zpupster):

sqroot should not be negative

OpenStudy (anonymous):

I need to prove it with limits

OpenStudy (xapproachesinfinity):

take the limit of that function at x=-3/2 from both sides! if the limit exist and have a value then the function must be continues past -1.5

OpenStudy (anonymous):

So I have to plug in numbers really close to -1.5 on both sides?

OpenStudy (xapproachesinfinity):

to show what im saying here \[\lim_{x \rightarrow \frac{-3^-}{2}}\sqrt{2x+3}=\lim_{x \rightarrow \frac{-3^+}{2}}\sqrt{2x+3}\]

OpenStudy (xapproachesinfinity):

we must show that this exist and true!

OpenStudy (xapproachesinfinity):

well that's one way of thinking about it^_^

OpenStudy (anonymous):

Would I say that the limit approaching from the left is undefined or DNE?

OpenStudy (xapproachesinfinity):

open study is going crazy lol

OpenStudy (anonymous):

what?

OpenStudy (xapproachesinfinity):

well check the limit and see! why would it be undefined?

OpenStudy (xapproachesinfinity):

to me it is tending to some value?

OpenStudy (anonymous):

Because the function doesn't exist before (-1.5,0)

OpenStudy (xapproachesinfinity):



OpenStudy (xapproachesinfinity):

but it is going towards some value even if there nothing before that point

OpenStudy (anonymous):

So I do not have to take the left hand limit because the function does not exist past -1.5?

OpenStudy (xapproachesinfinity):

yeah! the function needs not the part from the left!

OpenStudy (anonymous):

Ok. So I only take the right limit?

OpenStudy (xapproachesinfinity):

yes! your function will be continuos from x>=-3/2 other than that the function doesn't exist

OpenStudy (anonymous):

Ok. So I don't need to prove that it doesn't exist below that? I can just state that?

OpenStudy (xapproachesinfinity):

you could state the function is defined only in the interval x>=-3/2 domain of f and therefore limit from below doesn't exist

OpenStudy (anonymous):

Oh ok. Thanks!

OpenStudy (xapproachesinfinity):

yw!